OpenStudy (anonymous):

How do I find the secant 195 using the sum/difference identities?

OpenStudy (anonymous):

sec(195) = se(180+15) =- sec(15)=1.03

OpenStudy (anonymous):

Can you use the same identity you would use for cos?

OpenStudy (anonymous):

Sure. It can be easier if you express sec 195 in terms of cos. \[\sec 195^\text{o}=\frac{1}{\cos 195^\text{o}}\]

OpenStudy (anonymous):

Yes